at 16C two open end organ pipes when sounded together produce 51 beats in 3sec. How many beats per second will be produced if the temperature rises to 51C ?(Neglect the increase in length of pipe)

Grade:12

1 Answers

vishal kumar singh
7 years ago
since beat frequency(B)=difference of two individual frequecies.B=f1-f2.f=n×v/2L for open pipe.and we know,v=kT^1/2.B=(n1×k/2L-n2×k/2l)T^1\2now putting 51/3 and16+273=289 in above eqand in other eq with variable B and temp 51 + 273=324 and dividing both eq we will get B=18beats per second.cheers